OpenStudy (anonymous):

calculate the number of photons having a wavelength of 10.0 micrometers required to produce 1.0 kilojoules of energy

OpenStudy (anonymous):

Welcome to Openstudy

OpenStudy (anonymous):

To solve this problem, you can first calculate the energy of each photon by the formula \(\large E_{photon}=\frac{hc}{\lambda}\) then find the number of photon by \(\large N=\frac {E}{E_{photon}}\). \(h\) is the Planck's constant, \(6.626\times 10^{-34}\) J.s \(c\) is the speed of light, \(3\times 10^8\) m/s \(\lambda\) is the wavelength, in meter